# Separating complexity classes of LCL problems on grids

### v1

@@ 6.3.  Acknowledgments
The fourth named author is grateful to Riley Thornton for suggesting Proposition 2.18 as a possible tool for tackling Question 1.22. We are also grateful to Jan Grebík for insightful conversations, in particular for directing our attention to Question 1.17. Finally, we thank the organizers of the Workshop on Measurable Combinatorics that was held at the Alfréd Rényi Institute of Mathematics in Budapest, Hungary in June 2024, where part of this research took place.
This material is based upon work partially supported by the National Science Foundation under grants DGE-2146752, DMS-2239187, and DMS-2402064. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.

### v4

@@ 6.3.  Acknowledgments
The fourth named author is grateful to Riley Thornton for suggesting Proposition 2.18 as a possible tool for tackling Question 1.22, and to Andrew Marks for telling us about the source [Oberwolfach]. We are also grateful to Jan Grebík for insightful conversations, in particular for directing our attention to Question 1.17, and to Alexander Kechris and Anush Tserunyan for helpful comments on an earlier version of the manuscript. Finally, we thank the organizers of the Workshop on Measurable Combinatorics that was held at the Alfréd Rényi Institute of Mathematics in Budapest, Hungary in June 2024, where part of this research took place.
This material is based upon work partially supported by the National Science Foundation under grants DGE-2146752, DMS-2528522, and DMS-2402064. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.